# Attractive Metadata Attack: Inducing LLM Agents to Invoke Malicious Tools

### v1

@@ 3.3.3 Optimization Algorithm
TS ($\uparrow$) ASR ($\uparrow$) PR ($\uparrow$) PL ($\uparrow$) TS ($\uparrow$) ASR ($\uparrow$) PR ($\uparrow$) PL ($\uparrow$)
Gemma3-27B Injected Attack None 85.4085.4085.40 85.4085.4085.40 85.4085.4085.40 85.4085.4085.40 - - - -
Gemma3-27B Prompt Attack None 89.2089.2089.20 83.6083.6083.60 83.6083.6083.60 83.6083.6083.60 96.2096.2096.20 73.8073.8073.80 73.2073.2073.20 73.2073.2073.20
Gemma3-27B Our None 98.42 95.58 94.83 94.69 99.3099.3099.30 83.1083.1083.10 81.8081.8081.80 81.4981.4981.49
Gemma3-27B Our + Injected Attack None 95.3395.3395.33 95.3395.3395.33 94.5094.5094.50 94.1394.1394.13 99.60 99.20 98.20 97.61
Gemma3-27B Injected Attack Rewrite 80.6080.6080.60 78.0078.0078.00 (7.4)(-7.4)( - 7.4 ) 77.8077.8077.80 77.5177.5177.51 - - - -
Gemma3-27B Our Rewrite 95.3395.3395.33 90.5090.5090.50 (5.1)(-5.1)( - 5.1 ) 90.1290.1290.12 89.6589.6589.65 97.0097.0097.00 83.6083.6083.60 (+0.5)(+0.5)( + 0.5 ) 81.7481.7481.74 81.1981.1981.19
Gemma3-27B Our + Injected Attack Rewrite 91.8391.8391.83 91.0091.0091.00 (4.3)(-4.3)( - 4.3 ) 90.1790.1790.17 90.1790.1790.17 98.2098.2098.20 93.4093.4093.40 (5.8)(-5.8)( - 5.8 ) 91.6091.6091.60 91.2791.2791.27
Gemma3-27B Prompt Attack Refuge 96.0096.0096.00 84.6784.6784.67 (+1.1)(+1.1)( + 1.1 ) 83.5083.5083.50 83.3583.3583.35 92.3392.3392.33 53.0053.0053.00 (20.8)(-20.8)( - 20.8 ) 53.0053.0053.00 53.0053.0053.00
Gemma3-27B Our Refuge 96.0096.0096.00 89.0089.0089.00 (6.6)(-6.6)( - 6.6 ) 88.0088.0088.00 88.0088.0088.00 96.0096.0096.00 60.8060.8060.80 (22.3)(-22.3)( - 22.3 ) 59.2059.2059.20 58.4758.4758.47
Gemma3-27B Our + Injected Attack Refuge 97.33 97.33 (+2.0)(+2.0)( + 2.0 ) 94.67 94.67 100.00 100.00 (+0.8)(+0.8)( + 0.8 ) 97.20 96.61
Gemma3-27B Our + Injected Attack Rewrite + Refuge 94.3394.3394.33 94.3394.3394.33 (1.0)(-1.0)( - 1.0 ) 93.0093.0093.00 93.0093.0093.00 98.4098.4098.40 96.4096.4096.40 (2.8)(-2.8)( - 2.8 ) 94.8094.8094.80 93.6893.6893.68
LLaMA3.3-70B Injected Attack None 75.8075.8075.80 75.8075.8075.80 75.2075.2075.20 71.0471.0471.04 - - - -
LLaMA3.3-70B Prompt Attack None 99.2099.2099.20 90.4090.4090.40 90.4090.4090.40 90.4090.4090.40 97.2597.2597.25 74.0074.0074.00 73.5073.5073.50 73.5073.5073.50
LLaMA3.3-70B Our None 99.67 94.8094.8094.80 94.8094.8094.80 94.8094.8094.80 98.7598.7598.75 76.5576.5576.55 76.4876.4876.48 76.4576.4576.45
LLaMA3.3-70B Our + Injected Attack None 99.4799.4799.47 99.47 99.42 99.30 99.64 99.55 99.29 98.59
LLaMA3.3-70B Injected Attack Rewrite 87.4087.4087.40 70.0070.0070.00 (5.8)(-5.8)( - 5.8 ) 70.0070.0070.00 69.5669.5669.56 - - - -
LLaMA3.3-70B Our Rewrite 99.73 96.9396.9396.93 (+2.1)(+2.1)( + 2.1 ) 96.8096.8096.80 96.8096.8096.80 99.60 81.3081.3081.30 (+4.8)(+4.8)( + 4.8 ) 79.8779.8779.87 79.6979.6979.69
LLaMA3.3-70B Our + Injected Attack Rewrite 99.2099.2099.20 99.0799.0799.07 (0.4)(-0.4)( - 0.4 ) 99.0099.0099.00 98.8798.8798.87 99.60 98.3098.3098.30 (1.3)(-1.3)( - 1.3 ) 97.7397.7397.73 97.5797.5797.57
LLaMA3.3-70B Prompt Attack Refuge 96.5096.5096.50 84.5084.5084.50 (5.9)(-5.9)( - 5.9 ) 84.0084.0084.00 84.0084.0084.00 98.0098.0098.00 55.3355.3355.33 (18.7)(-18.7)( - 18.7 ) 55.3355.3355.33 55.3355.3355.33
LLaMA3.3-70B Our Refuge 98.6798.6798.67 90.4090.4090.40 (4.4)(-4.4)( - 4.4 ) 90.4090.4090.40 90.4090.4090.40 97.6097.6097.60 57.6057.6057.60 (19.0)(-19.0)( - 19.0 ) 57.6057.6057.60 57.4157.4157.41
LLaMA3.3-70B Our + Injected Attack Refuge 99.4799.4799.47 99.47 (+0.0)(+0.0)( + 0.0 ) 99.47 99.47 99.2099.2099.20 99.20 (0.4)(-0.4)( - 0.4 ) 99.17 98.36
LLaMA3.3-70B Our + Injected Attack Rewrite + Refuge 98.4098.4098.40 97.8797.8797.87 (1.6)(-1.6)( - 1.6 ) 97.8097.8097.80 97.6497.6497.64 98.0098.0098.00 94.2094.2094.20 (5.4)(-5.4)( - 5.4 ) 93.6193.6193.61 93.4493.4493.44
Qwen2.5-32B Injected Attack None 92.2092.2092.20 92.2092.2092.20 92.2092.2092.20 91.0691.0691.06 - - - -
Qwen2.5-32B Prompt Attack None 97.2097.2097.20 86.4086.4086.40 85.6085.6085.60 85.6085.6085.60 80.6080.6080.60 25.6025.6025.60 23.8023.8023.80 23.7823.7823.78
Qwen2.5-32B Our None 97.0897.0897.08 94.5494.5494.54 92.6992.6992.69 92.6392.6392.63 85.1085.1085.10 38.9538.9538.95 38.5538.5538.55 38.5338.5338.53
Qwen2.5-32B Our + Injected Attack None 99.69 99.69 99.69 99.63 98.70 98.70 98.60 98.53
Qwen2.5-32B Injected Attack Rewrite 90.4090.4090.40 76.0076.0076.00 (16.2)(-16.2)( - 16.2 ) 76.0076.0076.00 75.9975.$9$
Qwen2.5-32B Our Rewrite 97.3897.3897.38 94.9294.9294.92 (+0.4)(+0.4)( + 0.4 ) 93.8593.8593.85 93.8293.8293.82 82.2082.2082.20 36.8036.8036.80 (2.2)(-2.2)( - 2.2 ) 36.5036.5036.50 36.4936.4936.49
Qwen2.5-32B Our + Injected Attack Rewrite 99.6999.6999.69 99.6999.6999.69 (+0.0)(+0.0)( + 0.0 ) 99.5199.5199.51 99.4899.4899.48 98.40 98.30 (0.4)(-0.4)( - 0.4 ) 97.1297.1297.12 97.0497.0497.04
Qwen2.5-32B Prompt Attack Refuge 96.4096.4096.40 86.4086.4086.40 (+0.0)(+0.0)( + 0.0 ) 85.2085.2085.20 85.1485.1485.14 79.6779.6779.67 19.0019.0019.00 (6.6)(-6.6)( - 6.6 ) 18.1718.1718.17 18.1218.1218.12
Qwen2.5-32B Our Refuge 96.6296.6296.62 94.4694.4694.46 (0.1)(-0.1)( - 0.1 ) 93.5493.5493.54 93.5493.5493.54 80.8080.8080.80 33.4033.4033.40 (5.6)(-5.6)( - 5.6 ) 33.0033.0033.00 32.9832.9832.98
Qwen2.5-32B Our + Injected Attack Refuge 100.00 100.00 (+0.3)(+0.3)( + 0.3 ) 100.00 99.88 98.40 98.0098.0098.00 (0.7)(-0.7)( - 0.7 ) 97.80 97.77
Qwen2.5-32B Our + Injected Attack Rewrite + Refuge 98.4698.4698.46 98.4698.4698.46 (1.2)(-1.2)( - 1.2 ) 98.4698.4698.46 98.4398.4398.43 97.8097.8097.80 94.2094.2094.20 (4.5)(-4.5)( - 4.5 ) 94.0094.0094.00 93.8193.8193.81
GPT-4o-mini Injected Attack None 89.0089.0089.00 89.0089.0089.00 85.3385.3385.33 84.2084.2084.20 - - - -
GPT-4o-mini Prompt Attack None 81.2081.2081.20 72.4072.4072.40 72.4072.4072.40 72.1972.1972.19 40.8340.8340.83 6.836.836.83 6.836.836.83 6.836.836.83
GPT-4o-mini Our None 85.8685.8685.86 81.4381.4381.43 81.1481.1481.14 81.1281.1281.12 46.3046.3046.30 23.1023.1023.10 23.0023.0023.00 22.9922.9922.99
GPT-4o-mini Our + Injected Attack None 97.57 97.57 97.57 97.44 97.30 97.20 97.10 96.85
GPT-4o-mini Injected Attack Rewrite 71.0071.0071.00 63.0063.0063.00 (26.0)(-26.0)( - 26.0 ) 63.0063.0063.00 63.0063.$0$
GPT-4o-mini Our Rewrite 86.1586.1586.15 81.1581.1581.15 (0.3)(-0.3)( - 0.3 ) 81.1581.1581.15 80.8880.8880.88 54.2954.2954.29 26.5726.5726.57 (+3.5)(+3.5)( + 3.5 ) 26.2926.2926.29 26.2926.2926.29
GPT-4o-mini Our + Injected Attack Rewrite 93.0893.0893.08 93.0893.0893.08 (4.5)(-4.5)( - 4.5 ) 92.6992.6992.69 92.6992.6992.69 92.5792.5792.57 91.1491.1491.14 (6.1)(-6.1)( - 6.1 ) 90.8690.8690.86 90.8690.8690.86
GPT-4o-mini Prompt Attack Refuge 74.3374.3374.33 63.0063.0063.00 (9.4)(-9.4)( - 9.4 ) 63.0063.0063.00 62.9662.9662.96 36.2036.2036.20 3.603.603.60 (3.2)(-3.2)( - 3.2 ) 3.603.603.60 3.583.583.58
GPT-4o-mini Our Refuge 84.6484.6484.64 77.1477.1477.14 (4.3)(-4.3)( - 4.3 ) 76.7976.7976.79 76.6576.6576.65 41.5041.5041.50 11.1711.1711.17 (11.9)(-11.9)( - 11.9 ) 11.0011.0011.00 10.9810.9810.98
GPT-4o-mini Our + Injected Attack Refuge 97.86 97.86 (+0.3)(+0.3)( + 0.3 ) 96.07 95.85 96.86 96.86 (0.3)(-0.3)( - 0.3 ) 96.86 96.61
GPT-4o-mini Our + Injected Attack Rewrite + Refuge 91.0791.0791.07 91.0791.0791.07 (6.5)(-6.5)( - 6.5 ) 90.7190.7190.71 90.7190.7190.71 95.1495.1495.14 93.4393.4393.43 (3.8)(-3.8)( - 3.8 ) 93.4393.4393.43 93.2793.2793.27
@@ 4.1 Experimental Setup
We adopt the ReAct think–act–observe paradigm [24], implemented through AgentBench [14], and the security-focused benchmark ASB [29]. Following the experimental settings established in prior work, we simulate agent workflows across ten real-world scenarios spanning IT operations, portfolio management, and other domains. Each workflow consists of subtasks grounded in domain-specific APIs, designed to emulate realistic agent behavior. Building on this, each agent is further configured with a synthetic user profile from the AI4Privacy corpus [2], containing 11 standardized personally identifiable information (PII) fields (e.g., name, address, phone number) explicitly marked as non-disclosable in the system prompt. This setup enables systematic evaluation of privacy leakage risks. We evaluate the effectiveness of AMA on four mainstream LLMs: three open-source models—Gemma-3 27B [19], LLaMA-3.3-Instruct 70B [9], and Qwen-2.5-Instruct 32B [23]—and one commercial model, GPT-4o-mini [16, 11]. Additional implementation details are provided in Appendix LABEL:compute_environment.
@@
We consider two types of adversarial threat settings based on the attacker’s knowledge of the task context. In targeted attacks, the adversary has detailed knowledge of the agent’s domain and available tools (e.g., portfolio_manager in finance, prescription_manager in healthcare). In contrast, untargeted attacks assume no such contextual or tool-specific information. For more details on the AMA optimization configuration, please refer to Appendix LABEL:optimization_configuration. Detailed ablation studies regarding declared tool parameters and tool generation efficiency are provided in Appendices LABEL:declared_tool_parameters and LABEL:efficiency.
@@
We compare AMA against two representative baseline attack strategies: Injected Attack [15], which overrides the agent’s intent by appending imperative instructions that coerce specific tool usage; and Prompt Attack, which leverages prompt engineering to steer the LLM into generating malicious tool metadata during tool creation. Moreover, we further evaluate the effectiveness of these attacks under two defense mechanisms: Dynamic Prompt Rewriting (Rewrite) [29] rewrites user queries to preserve the original intent and filter out injected content; and Prompt Refuge (Refuge) embeds rule-based security guardrails into the system prompt, instructing the agent to reject tools whose metadata or behavior appear adversarial or anomalous.
@@ 4.2 Main Results
AMA achieves high attack success while preserving task performance. Table 1 shows that AMA consistently outperforms the baseline Prompt Attack across all evaluation metrics, including both targeted and untargeted settings, and on all four models. In the targeted setting, AMA achieves up to 94% ASR and 92% PL on the open-source models (Gemma-3 27B, LLaMA-3.3 70B, Qwen-2.5 32B), surpassing Prompt Attack by 4–12%. Compared to Injected Attack, AMA also delivers higher ASR on all open-source models (a 2–19% gain), and performs only slightly worse on the commercial GPT-4o-mini, with a gap of just 7.8%. When combined with Injected Attack, AMA further boosts performance across most models, including GPT-4o-mini. This stacked attack combines imperative steering with metadata manipulation, making the malicious tool appear both safe and optimal. As a result, ASR and PL increase to 94–99% across the board. The only exception is Gemma-3 27B, which shows a minor drop of 0.25%, while all other models continue to improve.
Moreover, AMA has negligible impact on task completion. On open-source models, the task success rate (TS) remains as high as 98%, with only a slight drop on GPT-4o-mini, indicating strong stealth and stability. In the untargeted setting, AMA exhibits similar trends. Although its performance is slightly reduced compared to the targeted setting, the attack remains highly effective, demonstrating strong generalizability even without contextual knowledge. To further assess the robustness of AMA, we conduct additional experiments on a recently released, reasoning-capable open-source LLM—Qwen3-32B. For more detailed results and analysis, please refer to the Appendix LABEL:attack_defense_results_qwen3.
Prompt-level defenses fail against AMA. Existing prompt-level defenses—Rewrite and Refuge—exhibit limited and inconsistent mitigation against AMA. Overall, Rewrite proves largely ineffective and may even exacerbate the attack, while Refuge offers only preliminary protection. Specifically, Rewrite reduces ASR by 5%–26% against Injected Attack, consistent with findings reported in ASB. However, once AMA is introduced, both defenses fail to provide effective protection and are rendered nearly useless. Similar trends are observed across other evaluation metrics, further highlighting AMA’s robustness and ability to bypass prompt-level safeguards.

### v2

@@ 16  Select top $n_{t}$ tools from $CT_{k}$ based on highest $v_{i}^{j}$ scores;
TS (↑) ASR (↑) PR (↑) PL (↑) TS (↑) ASR (↑) PR (↑) PL (↑)
Gemma3-27B Injection Attack None $85.40$ $85.40$ $85.40$ $85.40$ - - - -
Gemma3-27B Prompt Attack None $89.20$ $83.60$ $83.60$ $83.60$ $96.20$ $73.80$ $73.20$ $73.20$
Gemma3-27B Our None 98.42 95.58 94.83 94.69 $99.30$ $83.10$ $81.80$ $81.49$
Gemma3-27B Our + Injection Attack None $95.33$ $95.33$ $94.50$ $94.13$ 99.60 99.20 98.20 97.61
Gemma3-27B Injection Attack Rewrite $80.60$ $78.00$ (7.4)(-7.4) $77.80$ $77.51$ - - - -
Gemma3-27B Our Rewrite $95.33$ $90.50$ (5.1)(-5.1) $90.12$ $89.65$ $97.00$ $83.60$ $(+0.5)$ $81.74$ $81.19$
Gemma3-27B Our + Injection Attack Rewrite $91.83$ $91.00$ (4.3)(-4.3) $90.17$ $90.17$ $98.20$ $93.40$ (5.8)(-5.8) $91.60$ $91.27$
Gemma3-27B Prompt Attack Refuge $96.00$ $84.67$ $(+1.1)$ $83.50$ $83.35$ $92.33$ $53.00$ (20.8)(-20.8) $53.00$ $53.00$
Gemma3-27B Our Refuge $96.00$ $89.00$ (6.6)(-6.6) $88.00$ $88.00$ $96.00$ $60.80$ (22.3)(-22.3) $59.20$ $58.47$
Gemma3-27B Our + Injection Attack Refuge 97.33 97.33 $(+2.0)$ 94.67 94.67 100.00 100.00 $(+0.8)$ 97.20 96.61
Gemma3-27B Our + Injection Attack Rewrite + Refuge $94.33$ $94.33$ (1.0)(-1.0) $93.00$ $93.00$ $98.40$ $96.40$ (2.8)(-2.8) $94.80$ $93.68$
LLaMA3.3-70B Injection Attack None $75.80$ $75.80$ $75.20$ $71.04$ - - - -
LLaMA3.3-70B Prompt Attack None $99.20$ $90.40$ $90.40$ $90.40$ $97.25$ $74.00$ $73.50$ $73.50$
LLaMA3.3-70B Our None 99.67 $94.80$ $94.80$ $94.80$ $98.75$ $76.55$ $76.48$ $76.45$
LLaMA3.3-70B Our + Injection Attack None $99.47$ 99.47 99.42 99.30 99.64 99.55 99.29 98.59
LLaMA3.3-70B Injection Attack Rewrite $87.40$ $70.00$ (5.8)(-5.8) $70.00$ $69.56$ - - - -
LLaMA3.3-70B Our Rewrite 99.73 $96.93$ $(+2.1)$ $96.80$ $96.80$ 99.60 $81.30$ $(+4.8)$ $79.87$ $79.69$
LLaMA3.3-70B Our + Injection Attack Rewrite $99.20$ $99.07$ (0.4)(-0.4) $99.00$ $98.87$ 99.60 $98.30$ (1.3)(-1.3) $97.73$ $97.57$
LLaMA3.3-70B Prompt Attack Refuge $96.50$ $84.50$ (5.9)(-5.9) $84.00$ $84.00$ $98.00$ $55.33$ (18.7)(-18.7) $55.33$ $55.33$
LLaMA3.3-70B Our Refuge $98.67$ $90.40$ (4.4)(-4.4) $90.40$ $90.40$ $97.60$ $57.60$ (19.0)(-19.0) $57.60$ $57.41$
LLaMA3.3-70B Our + Injection Attack Refuge $99.47$ 99.47 $(+0.0)$ 99.47 99.47 $99.20$ 99.20 (0.4)(-0.4) 99.17 98.36
LLaMA3.3-70B Our + Injection Attack Rewrite + Refuge $98.40$ $97.87$ (1.6)(-1.6) $97.80$ $97.64$ $98.00$ $94.20$ (5.4)(-5.4) $93.61$ $93.44$
Qwen2.5-32B Injection Attack None $92.20$ $92.20$ $92.20$ $91.06$ - - - -
Qwen2.5-32B Prompt Attack None $97.20$ $86.40$ $85.60$ $85.60$ $80.60$ $25.60$ $23.80$ $23.78$
Qwen2.5-32B Our None $97.08$ $94.54$ $92.69$ $92.63$ $85.10$ $38.95$ $38.55$ $38.53$
Qwen2.5-32B Our + Injection Attack None 99.69 99.69 99.69 99.63 98.70 98.70 98.60 98.53
Qwen2.5-32B Injection Attack Rewrite $90.40$ $76.00$ (16.2)(-16.2) $76.00$ $75.99$ - - - -
Qwen2.5-32B Our Rewrite $97.38$ $94.92$ $(+0.4)$ $93.85$ $93.82$ $82.20$ $36.80$ (2.2)(-2.2) $36.50$ $36.49$
Qwen2.5-32B Our + Injection Attack Rewrite $99.69$ $99.69$ $(+0.0)$ $99.51$ $99.48$ 98.40 98.30 (0.4)(-0.4) $97.12$ $97.04$
Qwen2.5-32B Prompt Attack Refuge $96.40$ $86.40$ $(+0.0)$ $85.20$ $85.14$ $79.67$ $19.00$ (6.6)(-6.6) $18.17$ $18.12$
Qwen2.5-32B Our Refuge $96.62$ $94.46$ (0.1)(-0.1) $93.54$ $93.54$ $80.80$ $33.40$ (5.6)(-5.6) $33.00$ $32.98$
Qwen2.5-32B Our + Injection Attack Refuge 100.00 100.00 $(+0.3)$ 100.00 99.88 98.40 $98.00$ (0.7)(-0.7) 97.80 97.77
Qwen2.5-32B Our + Injection Attack Rewrite + Refuge $98.46$ $98.46$ (1.2)(-1.2) $98.46$ $98.43$ $97.80$ $94.20$ (4.5)(-4.5) $94.00$ $93.81$
GPT-4o-mini Injection Attack None $89.00$ $89.00$ $85.33$ $84.20$ - - - -
GPT-4o-mini Prompt Attack None $81.20$ $72.40$ $72.40$ $72.19$ $40.83$ $6.83$ $6.83$ $6.83$
GPT-4o-mini Our None $85.86$ $81.43$ $81.14$ $81.12$ $46.30$ $23.10$ $23.00$ $22.99$
GPT-4o-mini Our + Injection Attack None 97.57 97.57 97.57 97.44 97.30 97.20 97.10 96.85
GPT-4o-mini Injection Attack Rewrite $71.00$ $63.00$ (26.0)(-26.0) $63.00$ $63.00$ - - - -
GPT-4o-mini Our Rewrite $86.15$ $81.15$ (0.3)(-0.3) $81.15$ $80.88$ $54.29$ $26.57$ $(+3.5)$ $26.29$ $26.29$
GPT-4o-mini Our + Injection Attack Rewrite $93.08$ $93.08$ (4.5)(-4.5) $92.69$ $92.69$ $92.57$ $91.14$ (6.1)(-6.1) $90.86$ $90.86$
GPT-4o-mini Prompt Attack Refuge $74.33$ $63.00$ (9.4)(-9.4) $63.00$ $62.96$ $36.20$ $3.60$ (3.2)(-3.2) $3.60$ $3.58$
GPT-4o-mini Our Refuge $84.64$ $77.14$ (4.3)(-4.3) $76.79$ $76.65$ $41.50$ $11.17$ (11.9)(-11.9) $11.00$ $10.98$
GPT-4o-mini Our + Injection Attack Refuge 97.86 97.86 $(+0.3)$ 96.07 95.85 96.86 96.86 (0.3)(-0.3) 96.86 96.61
GPT-4o-mini Our + Injection Attack Rewrite + Refuge $91.07$ $91.07$ (6.5)(-6.5) $90.71$ $90.71$ $95.14$ $93.43$ (3.8)(-3.8) $93.43$ $93.27$
@@ 4.1 Experimental Setup
We adopt the ReAct think-act-observe paradigm Yao et al. (2023), implemented through AgentBench Liu et al. (2023), and the security-focused benchmark ASB Zhang et al. (2025). Following the experimental settings established in prior work, we simulate agent workflows across ten real-world scenarios spanning IT operations, portfolio management, and other domains. Each workflow consists of subtasks grounded in domain-specific APIs, designed to emulate realistic agent behavior. Building on this setup, each agent is further configured with a synthetic user profile from the AI4Privacy corpus Ai4Privacy (2025), containing 11 standardized personally identifiable information (PII) fields (e.g., name, address, phone number) explicitly marked as non-disclosable in the system prompt. This setup enables systematic evaluation of privacy leakage risks. We evaluate the effectiveness of AMA on four mainstream LLMs: three open-source models-Gemma-3 27B Team et al. (2025), LLaMA-3.3-Instruct 70B Grattafiori et al. (2024), and Qwen-2.5-Instruct 32B Yang et al. (2024)-and one commercial model, GPT-4o-mini OpenAI (2022); Hurst et al. (2024). Additional implementation details are provided in Appendix B.1.1.
@@
We consider two types of adversarial threat settings based on the attacker’s knowledge of the task context. In targeted attacks, the adversary has detailed knowledge of the agent’s domain and available tools (e.g., portfolio_manager in finance, prescription_manager in healthcare). In contrast, untargeted attacks assume no such contextual or tool-specific information. For more details on the AMA optimization configuration, please refer to Appendix B.1.2. Detailed ablation studies regarding declared tool parameters and tool generation efficiency are provided in Appendices B.2.1 and B.2.2.
@@
We compare AMA against two representative baseline attack strategies: Injection Attack Liu et al. (2024), which overrides the agent’s intent by appending imperative instructions that coerce specific tool usage; and Prompt Attack, which leverages prompt engineering to steer the LLM into generating malicious tool metadata during tool creation. Moreover, we further evaluate the effectiveness of these attacks under two defense mechanisms: Dynamic Prompt Rewriting (Rewrite) Zhang et al. (2025) rewrites user queries to preserve the original intent and filter out injected content; and Prompt Refuge (Refuge) embeds rule-based security guardrails into the system prompt, instructing the agent to reject tools whose metadata or behavior appear adversarial or anomalous.
@@ 4.2 Main Results
AMA achieves high attack success while preserving task performance. Table 1 shows that AMA consistently outperforms the baseline Prompt Attack across all evaluation metrics, including both targeted and untargeted settings, and on all four models. In the targeted setting, AMA achieves up to 94% ASR and 92% PL on the open-source models (Gemma-3 27B, LLaMA-3.3 70B, Qwen-2.5 32B), surpassing Prompt Attack by 4-12%. Compared to Injection Attack, AMA also delivers higher ASR on all open-source models (a 2-19% gain), and performs only slightly worse on the commercial GPT-4o-mini, with a gap of just 7.8%. When combined with Injection Attack, AMA further boosts performance across most models, including GPT-4o-mini. This stacked attack combines imperative steering with metadata manipulation, making the malicious tool appear both safe and optimal. As a result, ASR and PL increase to 94-99% across the board. The only exception is Gemma-3 27B, which shows a minor drop of 0.25%, while all other models continue to improve.
Moreover, AMA exerts negligible impact on task completion. On open-source models, the task success rate (TS) remains as high as 98%, with only a slight drop on GPT-4o-mini, indicating strong stealth and stability. In the untargeted setting, AMA exhibits similar trends. Although its performance is slightly reduced compared to the targeted setting, the attack remains highly effective, demonstrating strong generalizability even without contextual knowledge. To further assess the robustness of AMA, we conduct additional experiments on a recently released, reasoning-capable open-source LLM-Qwen3-32B. For more detailed results and analysis, please refer to the Appendix B.2.3.
Prompt-level defenses fail against AMA. Existing prompt-level defenses—Rewrite and Refuge—exhibit limited and inconsistent mitigation against AMA. Overall, Rewrite proves largely ineffective and may even exacerbate the attack, while Refuge offers only preliminary protection. Specifically, Rewrite reduces ASR by 5%-26% against Injection Attack, consistent with findings reported in ASB. However, once AMA is introduced, both defenses fail to provide effective protection and are rendered nearly useless. Similar trends are observed across other evaluation metrics, further highlighting AMA’s robustness and ability to bypass prompt-level safeguards.
Auditor-based detection struggles to detect AMA effectively. We further examine whether an additional “tool auditor” signal is capable of detecting AMA under both targeted and untargeted settings Jia et al. (2025). Specifically, we provide the GPT-4o-based auditor with sets of $\geq$20 tools, each containing one malicious tool generated by AMA, Prompt Attack, or Injection Attack along with several benign tools, and instruct it to flag suspicious tools based on their metadata. We repeat the auditing process ten times on varying tool sets and report the average results. As shown in Table 2, AMA is easier to detect than the Prompt Attack. Nevertheless, the detection rate of AMA by the auditor remains considerably lower than that of the Injection Attack (TPR = 55.00%), indicating its superior stealth. Moreover, AMA induces abnormal agent behavior more effectively, achieving a desirable trade-off between attack effectiveness and stealthiness.
@@
Table 2: Auditor-based detection performance under targeted and untargeted settings. We report TPR (malicious tools correctly detected) and FPR (benign tools incorrectly flagged). “-” indicates not applicable.
LLM Attack Setting Targeted Untargeted
TPR (↑) FPR (↓) TPR (↑) FPR (↓)
gemma-3-it Our 1.67 0.44 0.50 0.10
gemma-3-it Prompt Attack 1.50 0.39 0.43 0.18
llama-3.3-instruct Our 5.56 0.29 2.09 0.09
llama-3.3-instruct Prompt Attack 4.50 0.28 1.22 0.18
qwen2.5-instruct-32b Our 2.90 0.21 6.25 0.26
qwen2.5-instruct-32b Prompt Attack 2.32 0.22 4.20 0.43
gpt-4o-mini Our 2.39 0.44 2.10 0.35
gpt-4o-mini Prompt Attack 1.97 0.71 1.67 0.70
- Injection Attack 55.00 0.00 - -